# Variational Bayesian latent estimation with compressive autoencoders

### v2

@@ Abstract
Regularization of inverse problems is of paramount importance in computational imaging. The ability of neural networks to learn efficient image representations has been recently exploited to design powerful data-driven regularizers. While state-of-the-art plug-and-play methods rely on an implicit regularization provided by neural denoisers, alternative Bayesian approaches consider Maximum A Posteriori (MAP) estimation in the latent space of a generative model, thus with an explicit regularization. However, state-of-the-art deep generative models require a huge amount of training data compared to denoisers. Besides, their complexity hampers the optimization involved in latent MAP derivation. In this work, we first propose to use compressive autoencoders instead. These networks, which can be seen as variational autoencoders with a flexible latent prior, are smaller and easier to train than state-of-the-art generative models. As a second contribution, we introduce the Variational Bayes Latent Estimation (VBLE) algorithm, which performs latent estimation within the framework of variational inference. Thanks to a simple yet efficient parameterization of the variational posterior, VBLE allows for fast and easy (approximate) posterior sampling. Experimental results on image datasets BSD and FFHQ demonstrate that VBLE reaches similar performance than state-of-the-art plug-and-play methods, while being able to quantify uncertainties faster than other existing posterior sampling techniques.
@@ 1 Introduction
Image restoration tasks, such as deblurring, inpainting or super-resolution, consist in recovering a clean image $x$ from its noisy measurement $y$, based on the forward model $y=Ax+w$, with $A$ representing the degradation operator and $w$ an additive noise term. Within the Bayesian framework, this ill-posed inverse problem is typically solved by finding the Maximum A Posteriori (MAP) estimate, solution to the problem
@@
where $\log p_{Y|X}(y|x)$ is the observation log-likelihood, which is quadratic for Gaussian noise, and $-\log p_{X}(x)$ acts as a regularization term, promoting solutions that are most compatible with the prior distribution $p_{X}(x)$. Classical regularizations include total variation [1], Tikhonov regularization [2], and sparsity-promoting penalties on well-chosen representations such as wavelet bases or dictionaries [3, 4, 5].
Deep learning has led to substantial performance gains in image restoration tasks. A first category of methods directly solves the inverse problem after a supervised end-to-end training on a dataset of original-degraded image pairs related by the forward model [6, 7]. Their performance are impressive but only for the specific inverse problem considered during training. For that reason, a second category of methods aims at learning the regularization only, leveraging the forward model to solve the inverse problem. Besides offering a better interpretability, they allow to solve a wide range of inverse problems with the same neural network.
Among them, plug-and-play (PnP) image restoration methods [8] yield excellent performance on a wide variety of image restoration tasks. These methods classically use splitting algorithms, such as ADMM [8, 9] or HQS [10, 11], that separately handle the data term and the regularization term in the optimization problem (1). Their key concept is to use Gaussian denoisers [12], and in particular deep denoisers [13] in place of the proximal operator of the regularization. More recently, the introduction of generative denoisers, such as denoising diffusion models [14, 15], within the PnP framework, has demonstrated remarkable performance in solving inverse problems [16, 17, 18].
@@
where $z$ is the latent variable to be optimized, $p_{Z}(z)$ is the generative latent prior and $p_{Y|Z}(y|z)=p_{Y|X}(y|G(z))$, that is $G$ is considered as a deterministic transformation from the latent to the image space. The original approach introduced in [19] is efficient for solving severely ill-posed inverse problems on images lying on a low-dimensional manifold, such as centered close-ups on human faces. Unfortunately, the manifold constraint $x=G(z)$ too strongly restricts the solution space when dealing with images lying on a high-dimensional manifold, which is the case for highly diverse and/or non-structured images, or when solving less severely ill-posed inverse problems. To address this limitation, [20, 21] permit small deviations from the generative manifold while [22, 23, 24] propose to jointly optimize $x$ and $z$ using splitting algorithms. Solving eq. 2 using Normalizing Flows (NF) [25] is also an option as it ensures the accessibility of any element in the image space [26, 27]. Even so, the quality of the solution remains highly dependent on the generative model. Unfortunately, the use of very deep state-of-the-art generative models makes the optimization of eq. 2 problematic. Indeed, the gradient descent often gets trapped in poor local minima [28] and it can be regarded as too computationally demanding for restoring a single image. Moreover, the training of very deep generative models require a lot of ideal images, which might not be accessible in some application contexts. In order to circumvent gradient descent, [22] uses the encoder of a variational autoencoder (VAE) as a stochastic approximate of the generative model posterior in an alternate optimization scheme on $(x,z)$, while [29] obtains promising performance by extending this approach to hierarchical VAEs [30], at the cost of additional image restoration hyperparameters.
@@
All the above mentioned methods only provide a single point estimate of the inverse problem solution. However, many applications require an estimation of the confidence behind this solution. To this end, several Bayesian methods enable to sample the posterior distribution $p_{X|Y}(x|y)$ of the inverse problem solution. Some methods yield stochastic solutions to the inverse problem, relying, for instance, on the implicit prior provided by Gaussian denoisers [31] or exploiting diffusion model properties [18]. However, posterior sampling is then computationally expensive since the inverse problem must be solved for each sample generation. In this regard, Markov Chain Monte Carlo (MCMC) can lead to more effective techniques. In particular, [32] proposes to use Unadjusted Langevin Algorithm (ULA) to solve imaging inverse problems. Building on this, PnP-ULA [33] approximates the log-likelihood gradient in ULA by employing a Gaussian denoiser in a PnP framework. NF-ULA[34] directly approximates this gradient using normalizing flows while [35] uses a MCMC sampling scheme in the latent space of a generative model. Many of these methods offer theoretical convergence guarantees towards the actual posterior distribution. However  it requires in practice many iterations which makes the restoration computationally heavy, despite recent attempts to speed up Markov chain convergence [36, 37]. Additionally, convergence is conditioned to the tuning of sensitive hyperparameters. Consequently, these methods often struggle to reach the state-of-the-art performance and computational efficiency achieved by their point estimate counterparts, although yielding valuable uncertainty quantification.
@@
Focusing on latent optimization methods, this paper addresses the issues arising when dealing with diverse or non-structured images. Moreover, taking benefit of the latent framework, the proposed approach offers efficient means for posterior sampling. Hence, our contribution is twofold. First, we propose compressive autoencoders (CAEs) [38] as an alternative to state-of-the-art generative models for latent optimization methods. We consider in particular CAEs with a hyperprior [39], which yield excellent results in compression and can be seen as VAEs with an adaptive latent prior. In addition, they are significantly smaller than state-of-the-art generative models. On one hand, their adaptability allows to properly regularize inverse problems on non-structured images. On the other hand, their light structure makes the latent optimization using gradient descent [19] both scalable and effective. Second, we introduce the Variational Bayes Latent Estimation (VBLE) algorithm which estimates the latent posterior $p_{Z|Y}(z|y)$ by variational inference, leveraging CAE characteristics to design a simple yet efficient approximation of $p_{Z|Y}(z|y)$. In this way, VBLE enables to estimate the posterior distribution with negligible additional computational cost.
@@
We conduct a comprehensive set of experiments on FFHQ [40] and BSD [41] datasets, comprising three different inverse problems: deblurring, single image super resolution (SISR) and inpainting. The proposed approach yields competitive image restoration results compared to state-of-the-art methods. It additionally offers relevant posterior sampling abilities compared to MCMC methods, with an algorithm being significantly faster.
@@ 2.1 Variational autoencoders
where KL denotes the Kullblack-Leibler divergence, measuring the distance between distributions. Hence, by maximizing the ELBO on a dataset $\mathcal{D}$ containing several images $x$, the log-likelihood of $\mathcal{D}$ is maximized, while the distance between the approximate and true posterior distributions is minimized. However, this version of the ELBO is intractable and can be rewritten as follows:
@@ 2.2 Variational compressive autoencoders
Machine learning and especially deep learning have raised a significant interest in the lossy image compression community. In the widespread transform coding framework, the image is transformed, the obtained image representation is quantized and subsequently compressed using a lossless entropy encoder [44]. The more decorrelated and sparse the transform is, the more effective the compression becomes. Several reference model-based methods use the wavelet transform [45] while state-of-the-art deep learning methods use data-dependent transforms provided by the encoder part of a compressive autoencoder (CAE) [38, 39, 46, 47]. CAEs are learned by minimizing a rate-distortion trade-off = $\mathcal{L}=\mbox{ Rate}+\alpha\times\mbox{Distortion }$ [38, 39]. On one hand, the quantization introduces distortion between the original and the decompressed image. This distortion is typically measured by the Mean Squared Error (MSE) between the network input and output. On the other hand, the entropy encoder design assumes a prior distribution or so-called entropy model $p$ for the quantized latent representation, denoted by $z$ in the following. However, $z$ actual distribution, denoted by $q$, generally differs from $p$. The rate is approximated by the Shannon cross entropy: Rate =$\mbox{Rate }=\mathbb{E}_{z\sim q}[-\log_{2}p(z)]$. This value is minimized when $p=q$, that is when the rate corresponds to the information entropy.
A specificity of CAE training comes from the fact that quantization is non-differentiable. It is thus approximated during training by the addition of a uniform noise, which classically models quantization noise [48]. This noise is akin to the noise introduced during the training of a VAE in the generative framework. In the following, we denote by $\bar{z}$ the representation before the quantization. The latent variable $z$ is thus a noisy version of $\bar{z}$. Furthermore, state-of-the-art CAEs incorporate a hyperprior [39, 46, 49]. This additional autoencoder takes $\bar{z}$ as input and estimates $z$’s mean $\mu^{z}$ and standard deviation $\sigma^{z}$, as illustrated in Fig. 1. The prior on $z$ is then defined as the $\bar{z}$-dependent factorized distribution $\prod_{k}\mathcal{N}(\mu^{z}_{k},(\sigma^{z}_{k})^{2})$ convolved by a uniform distribution to align with quantization. The hyperprior introduces another latent variable $h$, of much smaller dimension than $z$. $h$ is also quantized, entropy coded and transmitted.
@@
$p_{\theta}(x,z,h)$ $=p_{\theta}(x|z)p_{\theta}(z|h)p_{\theta}(h)\mbox{ with }p_{\theta}(x|z)=\prod_{k}\mathcal{N}\big{(}x_{k};\,D_{\theta}(z)_{k},\frac{1}{2\alpha\log(2)}\big{)},$
$p_{\theta}(z|h)$ $=\prod_{k}\Big{[}\mathcal{N}(z_{k};\,\mu^{z}_{k},(\sigma^{z}_{k})^{2})*U\big{(}z_{k};\,[-\frac{1}{2},\frac{1}{2}]\big{)}\Big{]}\text{ and }p_{\theta}(h)=\prod_{k}p_{\psi}(h_{k})$ (7)
@@
$q_{\phi}(z,h|x)=q_{\phi}(z|x,h)q_{\phi}(h|x)\mbox{ with }q_{\phi}(z|x,h)=\prod_{k}\mathcal{U}(z_{k};\,[\bar{z}_{k}-\frac{1}{2},\bar{z}_{k}+\frac{1}{2}])$
and $\text{and }q_{\phi}(h|x)=\prod_{k}\mathcal{U}(h_{k};\,[\bar{h}_{k}-\frac{1}{2},\bar{h}_{k}+\frac{1}{2}]).$ (8)
@@
$\mathcal{L}(x)$ $=\mathbb{E}_{q_{\phi}(z,h|x)}\big{[}\log q_{\phi}(z,h|x)-\log p_{\theta}(x|z,h)-\log p_{\theta}(z,h)\big{]}$ (9)
@@
Furthermore, state-of-the-art CAEs combine a second autoencoder with an autoregressive part to design an improved hyperprior [46, 47]. In this case, CAEs cannot be entirely seen as VAEs, although their formulation stay close.
@@ 3.2 Variational Bayes Latent Estimation (VBLE)
We still assume that $p_{Y|Z}(y|z)=p_{Y|X}(y|D_{\theta}(z))$. Recall that Bora et al. [19] method to solve inverse problems yields a MAP estimate $\arg\max_{z}p_{Z|Y}(z|y)$ in the latent space. Unlike this deterministic approach resulting in a point estimate, we wish to approximate $p_{Z|Y}(z|y)$ through variational inference, leveraging VAEs and CAEs latent structure to propose a simple posterior parameterization. Hence, we present a stochastic version of eq. 10, termed Variational Bayes Latent Estimation (VBLE), which is illustrated in Fig. 3. It consists in performing variational inference with the following parametric families:
$E_{\bar{z},a}$ =$=\left\{q_{\bar{z},a}(z)\big{|}\bar{z},a\in\mathbb{R}^{C\times M\times N},a>0\right\}$
@@
Note that we detail both CAE and VAE frameworks, as VBLE can be applied with both networks. These parametric families are based on the same distribution than the inference encoder distribution $q_{\phi}(z|x)$, i.e. uniform for CAEs and Gaussian for VAEs. Their parameters are the mean latent representation $\bar{z}$, and an additional $a$ parameter with the same dimensions as $\bar{z}$, which models the uncertainty related to each coefficient of $z$.
Hence, employing a similar framework as the VAE training procedure described in Sec. 2.1, we maximize the ELBO:
$\arg\max_{\bar{z},a}\mathcal{L}_{\bar{z},a}=\arg\max_{z,a}\mathbb{E}_{q_{\bar{z},a}(z)}\big{[}\log p_{Y|Z}(y|z)+\log p_{\theta}(z)-\log q_{\bar{z},a}(z)\big{]}$ (11)
with $\log p_{\theta}(z)$ being the log-likelihood of the latent prior, that is in particular $\log\mathcal{N}(z;0,I)$ for a traditional VAE, and $\log(2)\mbox{ Rate}(z,E_{h,\phi}(\bar{z}))$ for a CAE with a hyperprior, as introduced in eq. 10. Note that $\log q_{\bar{z},a}(z)=-\sum_{k}\log a_{k}$ up to a constant in both Gaussian and uniform cases. We can then use the following reparameterization, for $z\sim q_{\bar{z},a}(z)$,
$z=\left\{\begin{array}[]{ll}\bar{z}+a\odot\epsilon\mbox{ with }\epsilon\sim\prod_{k}\mathcal{U}(\epsilon_{k};\,[-\frac{1}{2},\frac{1}{2}])&\mbox{ (CAE case)}\\ \bar{z}+a\odot\epsilon\mbox{ with }\epsilon\sim\mathcal{N}(0,I)&\mbox{ (VAE case)}\end{array}\right.$
where $\odot$ denotes the Hadamard, point-wise product between two tensors. Hence, the expectation in eq. 11 can be taken on $\epsilon$ using the reparametrization trick [42]. Therefore, a SGVB estimate of the ELBO can be derived, leading to algorithm 1.
@@
$\bar{z}_{0}\in\mathbb{R}^{C\times M\times N}$, $a_{0}\in\mathbb{R}^{C\times M\times N}=(1)_{i,j,l}$, $k=0$, $\eta>0$
@@
Both MMSE-x and MMSE-z estimates are relevant and can be used in practice.

### v4

# Variational Bayes image restoration with compressive autoencoders

@@ Abstract
Regularization of inverse problems is of paramount importance in computational imaging. The ability of neural networks to learn efficient image representations has been recently exploited to design powerful data-driven regularizers. While state-of-the-art plug-and-play (PnP) methods rely on an implicit regularization provided by neural denoisers, alternative Bayesian approaches consider Maximum A Posteriori (MAP) estimation in the latent space of a generative model, thus with an explicit regularization. However, state-of-the-art deep generative models require a huge amount of training data compared to denoisers. Besides, their complexity hampers the optimization involved in latent MAP derivation. In this work, we first propose to use compressive autoencoders instead. These networks, which can be seen as variational autoencoders with a flexible latent prior, are smaller and easier to train than state-of-the-art generative models. As a second contribution, we introduce the Variational Bayes Latent Estimation (VBLE) algorithm, which performs latent estimation within the framework of variational inference. Thanks to a simple yet efficient parameterization of the variational posterior, VBLE allows for fast and easy (approximate) posterior sampling. Experimental results on image datasets BSD and FFHQ demonstrate that VBLE reaches similar performance as state-of-the-art PnP methods, while being able to quantify uncertainties significantly faster than other existing posterior sampling techniques. The code associated to this paper is available in https://github.com/MaudBqrd/VBLE.
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
Index terms— Image restoration, Inverse problems, Deep regularization, Posterior sampling
@@ 1 Introduction
Image restoration tasks, such as deblurring, inpainting, or super-resolution, consist in recovering a clean image $x$ from its noisy measurement $y$, based on the forward model $y=Ax+w$, with $A$ representing the degradation operator and $w$ an additive noise term. Within the Bayesian framework, this ill-posed inverse problem is typically solved by finding the Maximum A Posteriori (MAP) estimate, solution to the problem
@@
where $\log p_{Y|X}(y|x)$ is the observation log-likelihood, which is quadratic for Gaussian noise, and $\log p_{X}(x)$ acts as a regularization term, promoting solutions that are most compatible with the prior distribution $p_{X}(x)$. Classical regularizations include total variation [57], Tikhonov regularization [64], and sparsity-promoting penalties on well-chosen representations such as wavelet bases or dictionaries [63, 30, 46].
Deep learning has led to substantial performance gains in image restoration tasks. A first category of methods directly solves the inverse problem after a supervised end-to-end training on a dataset of original-degraded image pairs related by the forward model [26, 58]. Their performance is impressive but only for the specific inverse problem considered during training. For that reason, a second category of methods aims at learning the regularization only, leveraging the forward model to solve the inverse problem. Besides offering better interpretability, they allow to solve a wide range of inverse problems with the same neural network.
Among them, plug-and-play (PnP) image restoration methods [66] yield excellent performance on a wide variety of image restoration tasks. These methods classically use splitting algorithms, such as ADMM [66, 12] or HQS [70, 38], that separately handle the data term and the regularization term in the optimization problem (1). Their key concept is to use Gaussian denoisers [40], and in particular deep denoisers [71] in place of the proximal operator of the regularization. More recently, the introduction of generative denoisers, such as denoising diffusion models [36, 51], within the PnP framework, has demonstrated remarkable performance in solving inverse problems [42, 17, 73, 11, 56, 32].
@@
where $z$ is the latent variable to be optimized, $p_{Z}(z)$ is the generative latent prior, and $p_{Y|Z}(y|z)=p_{Y|X}(y|G(z))$, that is $G$ is considered as a deterministic transformation from the latent to the image space. The original approach introduced in [7] is efficient for solving severely ill-posed inverse problems on images lying on a low-dimensional manifold, such as centered close-ups on human faces. Unfortunately, the manifold constraint $x=G(z)$ too strongly restricts the solution space when dealing with images lying on a high-dimensional manifold, which is the case for highly diverse and/or non-structured images, or when solving less severely ill-posed inverse problems. To address this limitation, [23, 22] permit small deviations from the generative manifold while [34, 27, 28] propose to jointly optimize $x$ and $z$ using splitting algorithms. Solving eq. 2 using Normalizing Flows (NF) [25] is also an option as it ensures the accessibility of any element in the image space [1, 52]. Even so, the quality of the solution remains highly dependent on the generative model. Unfortunately, the use of very deep state-of-the-art generative models makes the optimization of eq. 2 problematic. Indeed, the gradient descent often gets trapped in poor local minima [20] and it can be regarded as too computationally demanding for restoring a single image. Moreover, the training of very deep generative models requires a lot of ideal images, which might not be accessible in some application contexts. To circumvent gradient descent, [34] uses the encoder of a variational autoencoder (VAE) as a stochastic approximate of the generative model posterior in an alternate optimization scheme on $(x,z)$, while [55] obtains promising performance by extending this approach to hierarchical VAEs [60], at the cost of additional image restoration hyperparameters. Alternatively, PULSE [48] uses gradient descent over a high probability hyper-sphere in the latent space. This allows to reduce the blur induced by the quadratic latent regularization associated with the Gaussian latent prior.
@@
All the above mentioned methods only provide a single point estimate of the inverse problem solution. However, many applications require an estimation of the confidence behind this solution. To this end, several Bayesian methods enable to sample the posterior distribution $p_{X|Y}(x|y)$ of the inverse problem solution. Some methods yield stochastic solutions to the inverse problem, relying, for instance, on the implicit prior provided by Gaussian denoisers [39] or exploiting diffusion model properties [73][42]. However, posterior sampling is then computationally expensive since the inverse problem must be solved for each sample generation. In this regard, Markov Chain Monte Carlo (MCMC) can lead to more effective techniques. In particular, [29] proposes to use Unadjusted Langevin Algorithm (ULA) to solve imaging inverse problems. Building on this, PnP-ULA [44] approximates the log-likelihood gradient in ULA by employing a Gaussian denoiser in a PnP framework. NF-ULA[10] directly approximates this gradient using normalizing flows while [37] uses a MCMC sampling scheme in the latent space of a generative model. Many of these methods offer theoretical convergence guarantees towards the actual posterior distribution. However  it requires in practice many iterations which makes the restoration computationally heavy, despite recent attempts to speed up Markov chain convergence [18, 53]. Interestingly, variational inference (VI) [45] consists in explicitly approximating the posterior distribution by a parametric distribution. This allows for a fast sampling procedure. However, the existing VI deep-learning-based methods essentially learn, in a supervised manner, a neural network that fits the variational parameters to solve a given inverse problem [69, 6, 65].
@@
Focusing on latent optimization methods, this paper addresses the issues arising when dealing with diverse or non-structured images. Moreover, taking benefit of the latent framework, the proposed approach offers efficient means for posterior sampling. Hence, our contribution is twofold. First, we propose compressive autoencoders (CAEs) [3] as an alternative to state-of-the-art generative models for latent optimization methods. We consider in particular CAEs with a hyperprior [4], which yield excellent results in compression and can be seen as VAEs with an adaptive latent prior. In addition, they are significantly smaller than state-of-the-art generative models. On one hand, their adaptability allows them to properly regularize inverse problems on non-structured images. On the other hand, their light structure makes the latent optimization using gradient descent [7] both scalable and effective. Second, we introduce the Variational Bayes Latent Estimation (VBLE) algorithm which estimates the latent posterior $p_{Z|Y}(z|y)$ by variational inference, leveraging CAE characteristics to design a simple yet efficient approximation of $p_{Z|Y}(z|y)$. In this way, VBLE enables to estimate the posterior distribution with negligible additional computational cost.
@@
We conduct a comprehensive set of experiments on FFHQ [41] and BSD [47] datasets, comprising three different inverse problems: deblurring, single image super-resolution (SISR), and inpainting. Additionally, we perform a thorough study to demonstrate the relevance of VBLE posterior distribution. The proposed approach yields competitive image restoration results compared to state-of-the-art methods. Furthermore, it outperforms other posterior sampling methods in terms of computation time and GPU load, while yielding consistent abilities for posterior sampling and uncertainty quantification compared to state-of-the-art baselines.
@@ 2.1 Variational autoencoders
where KL denotes the Kullblack-Leibler divergence, measuring the distance between distributions. Hence, by maximizing the ELBO on a dataset $\mathcal{D}$ containing several images $x$, the log-likelihood of $\mathcal{D}$ is maximized, while the distance between the approximate and the true posterior distributions is minimized. However, this version of the ELBO is intractable and can be rewritten as follows:
@@ 2.2 Variational compressive autoencoders
[…]
A specificity of CAE training comes from the fact that quantization is non-differentiable. It is thus approximated during training by the addition of a uniform noise, which classically models quantization noise [8]. This noise is akin to the noise introduced during the training of a VAE in the generative framework. In the following, we denote by $\bar{z}$ the representation before the quantization. The latent variable $z$ is thus a noisy version of $\bar{z}$.
Furthermore, state-of-the-art CAEs incorporate a hyperprior [4, 50, 13]. This additional autoencoder takes $\bar{z}$ as input and estimates $z$’s mean $\mu^{z}$ and standard deviation $\sigma^{z}$, as illustrated in Fig. 1. The prior on $z$ is then defined as the $\bar{z}$-dependent factorized distribution $\prod_{k}\mathcal{N}(\mu^{z}_{k},(\sigma^{z}_{k})^{2})$ convolved by a uniform distribution to align with quantization. The hyperprior introduces another latent variable $h$, of a much smaller dimension than $z$. $h$ is also quantized, entropy coded, and transmitted.
@@
$p_{\theta}(x,z,h)$ $=p_{\theta}(x|z)p_{\theta}(z|h)p_{\theta}(h)$ (7)
$\mbox{with }p_{\theta}(x|z)$ $=\prod_{k}\mathcal{N}\big{(}x_{k};D_{\theta}(z)_{k},\frac{1}{2\alpha\log(2)}\big{)}$
$p_{\theta}(z|h)$ $=\prod_{k}\Big{[}\mathcal{N}(\mu^{z}_{k},(\sigma^{z}_{k})^{2})*U\big{(}z_{k};\,[-\frac{1}{2},\frac{1}{2}]\big{)}\Big{]}$
$p_{\theta}(h)$ $=\prod_{k}p_{\psi}(h_{k})$
@@
$q_{\phi}(z,h|x)$ $=q_{\phi}(z|x,h)q_{\phi}(h|x)$ (8)
$\mbox{with }q_{\phi}(z|x,h)$ $=\prod_{k}\mathcal{U}(z_{k};[\bar{z}_{k}-\frac{1}{2},\bar{z}_{k}+\frac{1}{2}]),$
$q_{\phi}(h|x)$ $=\prod_{k}\mathcal{U}(h_{k};[\bar{h}_{k}-\frac{1}{2},\bar{h}_{k}+\frac{1}{2}]).$
@@
$\mathcal{L}(x)$ $=\mathbb{E}_{q_{\phi}(z,h|x)}\big{[}\log q_{\phi}(z,h|x)-\log p_{\theta}(x|z,h)$
$\quad\quad\quad\quad\quad\quad-\log p_{\theta}(z,h)\big{]}$ (9)
@@
Furthermore, state-of-the-art CAEs combine a second autoencoder with an autoregressive part to design an improved hyperprior [50, 15]. In this case, CAEs cannot be entirely seen as VAEs, although their formulation stays close.
@@ 3.2 Variational Bayes Latent Estimation (VBLE)
We still assume that $p_{Y|Z}(y|z)=p_{Y|X}(y|D_{\theta}(z))$. Recall that Bora et al. [7] method to solve inverse problems yields a MAP estimate $\arg\max_{z}p_{Z|Y}(z|y)$ in the latent space. Unlike this deterministic approach resulting in a point estimate, we wish to approximate $p_{Z|Y}(z|y)$ through variational inference (VI), leveraging VAEs and CAEs latent structure to propose a simple posterior parameterization. Hence, we present a stochastic version of eq. 10, termed Variational Bayes Latent Estimation (VBLE), which is illustrated in Fig. 3. It consists of performing VI with the following parametric families:
${\mathcal{E}}_{\bar{z},a}$ $=\left\{q_{\bar{z},a}(z)\big{|}\bar{z},a\in\mathbb{R}^{C\times M\times N},a>0\right\}$ (11)
$\text{ with }q_{\bar{z},a}(z)$ $=\prod_{k}\mathcal{U}(z_{k};[\bar{z}_{k}-\frac{a_{k}}{2},\bar{z}_{k}+\frac{a_{k}}{2}])\quad\text{(CAE case),}$
$q_{\bar{z},a}(z)$ $=\prod_{k}\mathcal{N}(z_{k};\bar{z}_{k},a_{k}^{2})\quad\text{(VAE case).}$
Note that we detail both CAE and VAE frameworks, as VBLE can be applied with both of them. These parametric families are based on the same distribution as the inference encoder distribution $q_{\phi}(z|x)$, i.e. uniform for CAEs and Gaussian for VAEs. Their parameters are the mean latent representation $\bar{z}$, and an additional $a$ parameter with the same dimensions as $\bar{z}$, which models the uncertainty related to each coefficient of $z$. A representation of the posterior $q_{\bar{z},a}(z)$ is given in Fig. 2. The choice of a posterior belonging to the same distribution family as the encoder posterior $q_{\phi}(z|x)$ is a simple yet relevant choice. Indeed the encoder is trained so that sampling $q_{\phi}(z|x)$ provides the most likely latent variables given an input image $x$. Here, we aim to find the most likely latent variables given a measure $y$. Hence, adopting a unimodal distribution similar to the encoder posterior, however with an optimized variance, is a natural choice. Experiments in Sec. 4.4 demonstrate that this choice is beneficial to VBLE posterior sampling ability.
Then, employing a similar framework as the VAE training procedure described in Sec. 2.1, we maximize the ELBO:
$\arg\max_{\bar{z},a}\mathcal{L}_{\bar{z},a}$ $=\arg\max_{z,a}\mathbb{E}_{q_{\bar{z},a}(z)}\big{[}\log p_{Y|Z}(y|z)$
$+\log p_{\theta}(z)-\log q_{\bar{z},a}(z)\big{]}$ (12)
@@
where $\odot$ denotes the Hadamard, point-wise product between two tensors. Hence, the expectation in eq. 12 can be taken on $\epsilon$ using the reparametrization trick [43]. Therefore, an SGVB estimate of the ELBO can be derived, leading to algorithm 1.
@@
$y$ a degraded image, $\bar{z}_{0}\in\mathbb{R}^{C\times M\times N}$, $a_{0}\in\mathbb{R}^{C\times M\times N}=(1)_{i,j,l}$, $k=0$, $\eta>0$
@@
Both MMSE-x and MMSE-z estimates are relevant and can be used in practice. Note that this variational framework does not increase the restoration time compared to its deterministic counterpart MAP-z [7]. Indeed, after convergence of VBLE, sampling from the approximate posterior is fast since it only requires forward passes through the decoder, which can be parallelized within batches.